# Fine-grained Hallucination Detection and Editing for Language Models

### v3

@@ 5 Model: Fava
We introduce Fava (FAct Vericaton with Augmentation), a model for fine-grained hallucinations detections and editing. Fava is trained on high-quality synthetic training data to identify hallucinations, incorporating retrieved knowledge.
Overview. Fava consists of two components: a retriever $\mathcal{M}_{ret}$ and an editing LM $\mathcal{M}_{edit}$. $\mathcal{M}_{ret}$ takes the original output LM $y$ and optionally input prompt $x$ (if applicable), and retrieves top relevant documents: $\textbf{C}=\mathcal{M}_{ret}(x,y)$. Subsequently, the editing model detects and, if possible, edits factual errors in $y$ given the retrieved context: $\hat{y}=\mathcal{M}_{edit}(x,y,\textbf{C})$. $\hat{y}$ is an augmented output $y$ interleaved by the error edits with hallucination types as shown in Figure 1. While $\mathcal{M}_{edit}$ can be any LM, in our preliminary experiments, we find that making a state-of-the-art proprietary LM such as ChatGPT to perform fine-grained editing via prompting only is challenging (ref Table 2). Reliance on such API models also hurt reproducibility. Therefore, we generate high-quality synthetic training data with minimal human efforts (Section 5.1) and fine-tune a smaller yet powerful expert LM on the generated data (Section 5.2).
@@ 5.1 Synthetic Training Data Curation
To train our $\mathcal{M}_{edit}$ model, we require a large number of training instances, each of which includes $(c,y,y^{*})$, where $c$ is the gold context (original Wikipedia paragraph), $y$ is an erroneous LM output, and $y^{*}$ is an output with error tags and correct editing based on $c$. Fava leans to take $(c,y)$ and generate $y^{*}$. Yet, such data is not publicly available, and our annotation is expensive to scale up. Inspired by prior work that leverages LMs to generate synthetic training data Balachandran et al. (2022); Wang et al. (2023b); Asai et al. (2023), we introduce a new data creation method tailored to our fine-grained hallucination taxonomies.
Specifically, our data creation pipeline consists of three steps: Step 1: seed passage generations that generates a large, diverse collection of seed passages $c$, Step 2: error insertions which inserts factually incorrect statements informed by our our hallucination taxonomy to generate erronous passages $y$, and Step 3: post-processing which prunes and curate the final training examples. Figure 4 shows the overview of the pipeline.
@@
Seed passage generation. The goal in this step is to collect and generate a large, diverse range of seed passages to ensure Fava can handle a wide range of text types while maintaining control over factuality. Later on, errors are inserted into these seed passages to generate erroneous text, $y$. We base our seed passage collection on Wikipedia articles and available Question Answering (QA) datasets Kwiatkowski et al. (2019) since they are both good sources to evaluate factuality. However, the text genre is quite limited. To diversify the passage genres, we paraphrase Wikipedia passages into different genres to enable Fava to be adaptable and effective across various textual formats.
More specifically, we randomly sample 35,074 Wikipedia articles $c$ and use them as our gold reference passages. We then randomly sample one of the pre-specified genre types (e.g. blog article, tweet, etc.) for each passage and use ChatGPT to paraphrase the original passage into the given genre. In addition, we use 5,000 examples from QA datasets to further diversify the passage collections. In particular, we prompt ChatGPT with two QA demonstrations to generate 5,000 queries and correct answers using the information in the sampled Wikipedia articles. See the full list of the genres, instructions, and final ChatGPT-generated text in Appendix Table 7.
@@
Error insertion. The goal of this step is to simulate erroneous LM generation $y$, which are passages with factual errors. To control error distributions in the training data and incorporate our hallucination taxonomy, we prompt LMs to insert errors from our taxonomy one by one. In our pilot studies, we found that asking ChatGPT to insert multiple error types at the same time easily makes the model misunderstand or get confused with different error types (e.g., swapping or incorrectly identifying an error type). We also found while ChatGPT is capable of generating more simple types of errors such as  Entity , it struggles with generating plausible and difficult perturbations for more nuanced types. Therefore, we use ChatGPT and GPT-4 interchangeably for six different types. [^7] We use ChatGPT (gpt-3.5-turbo-0301) for the Type 1 and Type 3 while using GPT-4 for other types. Given an instruction and few-shot demonstrations of an error type $e^{type}$, GPT-4 or ChatGPT inserts new errors while retaining previously inserted errors and original text. The model was specifically instructed to mark phrases or sentences for deletion along with their error type and insert phrases and sentences along with insertion tags, allowing us to verify the error types and control their distribution. After this process, we have an erroneous text $y$, which inserts multiple errors into the original factual text $t$.
@@
Post-processing. We then post-process data, by swapping and removing the error tags and edits to form a clean erroneous text $y$ as the input and use the original correct passage $y^{*}$ as the output with edits, for training $\mathcal{M}_{edit}$. We filter out the examples violating our editing rules (e.g., inserting or removing errors without marking them up with tags and error types). We also retrieve additional four relevant Wikipedia paragraphs using Contriever-MSMARCO Izacard et al. (2022), and randomly mix the order of the references, forming the final references $\mathbf{C}$.
@@
Statistics of data.  After analyzing ablations with varying training instance sizes (Section 7.2), we generated 35,074 training instances,  [^8] While we see performance improvements up to 30k, the API costs for training data generation were already \$3k, reaching our initial budget. 30,0074 based on Wikipedia passages, and 5,000 based on QA pairs. All of the error types are almost equally distributed ($\sim$15 % each) with 3.1 errors inserted for Wikipedia subsets, and 1.4 errors inserted for the QA pairs on average. We present the error distributions in Appendix Table 8 and a detailed human evaluation in Appendix C.4.
@@ 5.2 Training and Inference
We train a smaller LM on the generated large-scale training data. In particular, given a training instance $(\mathbf{C},y,y^{*})$, our model $\mathcal{M}_{edit}$ takes $(\mathbf{C},y)$ as input and learns to predict the edited outputs with tags to represent error type $y^{*}$ using the standard language modeling objective. We use Llama2-Chat 7B to initialize $\mathcal{M}_{edit}$. We empirically found initializing Fava with this model gives slightly better performance than the pre-trained Llama2.
At inference time, we retrieve the top five documents from Wikipedia, [^9] We use English Wikipedia data from January 2023 and generate embeddings using the Contriever encoder. using Contriever-MSMARCO, and input them together with an LM output that may include hallucinations. The model identifies hallucinations, marks phrases or sentences for deletion, and suggests edits.
@@ 6 Experiments
Generator: ChatGPT Generator: Llama2-Chat 70B
Editor  ent  rel  con  inv  subj  unv OA Bi  ent  rel  con  inv  subj  unv OA Bi
ChatGPT 19.5 28.6 40.0 11.8 7.7 0.0 18.8 50.1 24.7 15.6 26.7 11.0 17.6 12.8 24.1 68.4
Rt+ChatGPT 28.1 19.2 25.5 5.4 37.7 15.5 24.4 64.8 33.7 24.2 24.0 22.2 17.8 4.7 27.8 72.8
GPT4 38.6 16.6 17.9 22.2 50.0 17.2 34.2 60.8 55.5 60.0 21.2 15.4 2.0 25.0 42.5 74.2
Fava (ours) 54.5 25.0 66.7 16.7 70.5 35.3 48.1 79.6 57.3 34.5 27.7 52.2 31.25 43.4 47.2 80.3
@@ 6.1 Experiments for Hallucination Detection
Evaluation data.  We use our new benchmark, consisting of 902 annotated passages, [^10] We collected 1010 annotated passages including our experimental batches, excluding earlier batches for evaluations. as our test data. We measure the models’ detection performance based on sentence-level per-category classification task as formulated in Section 3.2.
Baselines.  For fine-grained hallucination detection tasks, we test three baselines. ChatGPT prompts ChatGPT (gpt-3.5-turbo-0301) with a carefully designed prompt describing all six categories with two demonstrations. GPT4 (gpt-4) does the exact same but prompts GPT4 instead.We were unable to include GPT4 with retrieval in our baseline due to high costs. Since Fava uses retrieval augmentation, for comparability we include Rt+ChatGPT. This baseline prompts ChatGPT using the same prompt and demonstrations but also includes the top five retrieved documents by Contriever at test time to augment the original prompt. [^11] While we also tested strong white box LMs including Llama2-Chat 13B, we found their predictions often show confusion among different categories or struggle to output predictions in the expected format.
@@ 6.2 Experiments for Hallucination Editing
Evaluation data.  For editing evaluations, we use open-ended text generation data and evaluate models’ factuality based on a metric designed to measure the factuality of text, FActScore Min et al. (2023). This model-based metric prompts ChatGPT and GPT 3 (davinci-003) to decompose a response into a set of atomic facts and verify factuality for each using passages from a designated Wikipedia article. We use the original biography generation task proposed with FActScore.
@@ 6.3 Human Evaluations
Our automatic evaluations may not fully capture the models’ abilities to detect and refine hallucinations, due to the potential subjectivity of our annotations or the performance of the factuality evaluation metric. We evaluate randomly sampled 50 outputs from Fava as well as the baseline with the highest automatic evaluation score, namely Rt-ChatGPT. We ask human annotators to verify how many of the detection and edits are indeed correct based on the provided retrieved documents. This is similar to our automatic precision evaluation, but instead of coarsely evaluating detection performance at the sentence level, we evaluate the model performance at the individual detection level. Due to the cost of evaluating the factuality of all verification-worthy statements Min et al. (2023), we do not evaluate recall.
@@ 7.1 Results
Detection results.  Table 2 shows the fine-grained detection accuracy, namely F1 scores, and binary prediction F1 of Fava and baselines. We provide the full precision and recall scores in Appendix Tables 11 and 12, respectively. Fava significantly outperforms ChatGPT, Ret-ChatGPT, or GPT-4 on both fine-grained error detection and binary error detection. Fava shows high accuracy on error types such as   Sentence ,  Subjective ,  Entity . On the other hand, its performance on  Invented and  Unverifiable is still limited. Those two error types often require intensive search over many documents beyond the top few ones, while Fava by default only considers the top five documents.
@@
Editor ChatGPT Al-13B Al-7B
@@
Table 3: Overall Editing results on FActScore Min et al. (2023). Al-13B and Al-7B denote Alpaca 13B and 7B, respectively.
Editing results.  Table 3 shows the results of the editing task on FActScore. Our Fava significantly outperforms prompted ChatGPT or Llama2-Chat 70B, despite being a much smaller model in size. We found that ChatGPT is often confused with different errors and has a tendency to over-predict contradictory or subjective statements. Fava shows the largest gains in FActScore, showing the effectiveness of editing, in addition to error type detection.
7.2 Analysis
Effects of data scale.  We assess our model on varying sizes of synthetic training data (10k, 20k, and 30k instances) and examine their performance on fine-grained hallucination detection. Figure 4(b) demonstrates that Fava variants with a larger number of training instances perform significantly better at detecting fine-grained errors.
Effect of retrieval.  We assess Fava’s editing performance on Alpaca 13B outputs with varying retrieved evidence: top 1 document; top 5 documents; reordering the top 5 documents; top 4 documents and mix them with an introductory paragraph of the target entity (entity matching). Table 4(a) shows the experimental results, and more detailed results are available in Appendix Table  16.
We found that retrieving the top five documents significantly enhances performance compared to retrieval of only the top one document by 4.5%. Ordering passages using a reranking model gives some improvements, which indicates the non-negligible effect of context ordering, as reported by prior work Liu et al. (2023). Including the reference based on entity matching gives a notable enhancement to 50.1%, marking a 3.1% gain from the top five retrievals and 7.6% improvement over the unedited generations. These results indicate although Fava demonstrates strong capabilities, there’s room for improvement in retrieving and incorporating many references.
@@ 7.2 Analysis
(a) Retrieval
(b) Training Size
Figure 5: (a): FActScore results with different retrieval settings. (b): Detection scores across different training instance values.
@@ 8 Conclusions
Limitations
This work focuses on hallucinations in LM generations in an information-seeking context, where factual correct matters, and we build a new taxonomy, benchmark, and model for this. Errors in other contexts (e.g., commonsense reasoning, numerical reasoning) are beyond our focus. Our training data uses Wikipedia passages as gold context, and Fava uses Wikipedia passages as references at inference time. Yet, Wikipedia may not cover all of the necessary information to verify the factuality of LM-generated statements Asai and Choi (2021), and it is an open question how well Fava performs when it uses other types of documents as reference documents. Our analysis and findings have been based only on the English language and our scope is limited to information-seeking scenarios. Our model, Fava, and benchmark, FavaBench, were also trained on English text data. To understand fine-grained hallucinations in diverse languages or scenarios, more analysis and experiments need to be done.
Our annotations were collected with undergraduate and graduate computer science students who attended 45-minute training sessions to guarantee quality. Furthermore, annotations were reviewed for quality by us authors and annotator agreement was calculated to ensure understanding within annotators. Yet, given the complexity of our annotation tasks, there may be subjectivity or errors in annotations. We will release our data publicly to enable further investigations.
It is difficult to guarantee that our proposed hallucination taxonomy is all-encompassing. To address this concern, we asked annotators to include hallucination types not found in our taxonomy during the annotation process. We received no additions to our taxonomy during this, indicating our taxonomy is thorough.
We evaluate Fava to determine feasibility and quality. Although this doesn’t assure quality outputs from Fava consistently, the positive results indicate quality within our defined scope and setup.
Ethics Statement and Broader Impacts
LM-generated text is prone to hallucinations, which can result in severe real-world risks such as spreading misinformation. In this work, we attempt to understand such prevalent hallucinations by providing a new taxonomy, a human-annotated benchmark, and a new expert LM trained for fine-grained hallucinations. We hope our work will inspire future work and encourage more work focusing on fine-grained hallucination detections. Yet, our resulting models are not perfect and may overlook hallucinations in LM outputs and we acknowledge such limitations in our paper. For evaluating the detection capabilities of our model, we use our annotation data. Our annotation collection process did not involve gathering private information or data from our annotators. Additionally, of the four datasets we used for annotation prompts, three are publicly available datasets containing no personal information. The remaining prompt dataset was curated by us authors and does not include personal information. Our annotators were all paid for their time in completing annotations at the rate of \$20 per hour for 10 hours of work across a week. They all disclosed the payment amount and estimated time before beginning annotations. None of their data was collected while constructing the annotation dataset. For evaluating the editing capabilities of our model, we use FActScore Min et al. (2023) which has publicly available code and data and is based on Wikipedia.

### v4

@@ 5 Model: Fava
We introduce Fava (FAct Vericaton with Augmentation), a model for fine-grained hallucinations detections and editing. Fava is trained on high-quality synthetic training data to identify hallucinations, incorporating retrieved knowledge. Fava consists of two components: a retriever $\mathcal{M}_{ret}$ and an editing LM $\mathcal{M}_{edit}$. $\mathcal{M}_{ret}$ takes the original output LM $y$ and optionally input prompt $x$ (if applicable), and retrieves top relevant documents: $\textbf{C}=\mathcal{M}_{ret}(x,y)$. Subsequently, the editing model detects and, if possible, edits factual errors in $y$ given the retrieved context: $\hat{y}=\mathcal{M}_{edit}(x,y,\textbf{C})$. $\hat{y}$ is an augmented output $y$ interleaved by the error edits with hallucination types as shown in Figure 1. While $\mathcal{M}_{edit}$ can be any LM, in our preliminary experiments, we find that making a state-of-the-art proprietary LM such as ChatGPT to perform fine-grained editing via prompting only is challenging (Table 2). Reliance on such models also hurts reproducibility. Therefore, we generate high-quality synthetic training data with minimal human efforts (Section 5.1) and fine-tune a smaller yet powerful LM on the generated data (Section 5.2).
@@ 5.1 Synthetic Training Data Curation
To train our $\mathcal{M}_{edit}$ model, we require a large number of training instances, each of which includes $(c,y,y^{*})$, where $c$ is the gold context, $y$ is an erroneous LM output, and $y^{*}$ is an output with error tags and correct editing based on $c$. Fava learns to take $(c,y)$ and generate $y^{*}$. Inspired by prior work that leverages LMs to generate synthetic training data (Balachandran et al., 2022; Wang et al., 2023b; Asai et al., 2023), we introduce a new data creation method tailored to our fine-grained hallucination taxonomies. Specifically, our data creation pipeline (refer Figure 4) consists of three steps: Step 1: seed passage generations that generates a large, diverse collection of seed passages $c$, Step 2: error insertions which inserts factually incorrect statements to generate erronous passages $y$, and Step 3: post-processing which curates the final training examples.
@@
Seed passage generation. The goal in this step is to collect and generate a large, diverse range of seed passages to ensure Fava can handle a wide range of text types while maintaining control over factuality. Later on, errors are inserted into these seed passages to generate erroneous text, $y$. We base our seed passage collection on Wikipedia articles and sampled 5k QA pairs from Natural Questions (Kwiatkowski et al., 2019). We paraphrase Wikipedia passages into different genres to enable Fava to be adaptable and effective across various textual formats. Specifically, we randomly sample 35,074 Wikipedia articles $c$ and use them as our gold reference passages. We then randomly sample one of the pre-specified genre types (e.g. blog article, tweet, etc.) for each passage and use ChatGPT to paraphrase the original passage into the given genre. [^8] See the full list of the genres, instructions, and final ChatGPT-generated text in Appendix Table 9.
@@
Error insertion. The goal of this step is to simulate erroneous LM generation $y$, which are passages with factual errors. To control error distributions in the training data and incorporate our hallucination taxonomy, we prompt LMs to insert errors from our taxonomy one by one. We use ChatGPT and GPT-4 interchangeably for six different types. [^9] We use ChatGPT (gpt-3.5-turbo-0301) for the Type 1 and Type 3 while using GPT-4 for other types. See detailed discussions in Appendix B. Given an instruction and few-shot demonstrations of an error type $e^{type}$, GPT-4 or ChatGPT inserts new errors while retaining previously inserted errors. Models mark phrases or sentences for deletion along with their error type and insert phrases and sentences with insertion tags. Now we have an erroneous text $y$, which inserts multiple errors into the original text $t$.
@@
Post-processing. We then post-process data, by swapping and removing the error tags and edits to form a clean erroneous text $y$ as the input and use the original correct passage $y^{*}$ as the output with edits, for training $\mathcal{M}_{edit}$. We filter out the examples violating our editing rules (e.g., inserting errors without marking them up with tags and error types). We also retrieve additional four relevant Wikipedia paragraphs using Contriever-MSMARCO (Izacard et al., 2022), and randomly mix the order of the references, forming the final references $\mathbf{C}$.
@@
Statistics of data and human evaluation.  After analyzing ablations with varying training instance sizes (Section 7.2), we generated 35,074 training instances,  [^10] While we see performance improvements up to 30k, the API costs for training data generation were already \$3k, reaching our initial budget. 30,0074 based on Wikipedia passages, and 5,000 based on QA pairs. All of the error types are almost equally distributed ($\sim$15 % each) with 3.1 errors inserted for Wikipedia subsets, and 1.4 errors inserted for the QA pairs on average. We present the error distributions in Appendix Table 10. We conduct human evaluations on randomly sampled 50 examples and find that the generated data meets the expected criterion in most cases. See details in Appendix C.4.
@@ 5.2 Training and Inference
We train a smaller LM on the generated large-scale training data. In particular, given a training instance $(\mathbf{C},y,y^{*})$, our model $\mathcal{M}_{edit}$ takes $(\mathbf{C},y)$ as input and learns to predict the edited outputs with tags to represent error type $y^{*}$ using the standard language modeling objective. We use Llama2-Chat 7B to initialize $\mathcal{M}_{edit}$. We empirically found initializing Fava with this model gives slightly better performance than the pre-trained Llama2. At inference time, we retrieve the top five documents from Wikipedia, [^11] We use English Wikipedia data from January 2023 and generate embeddings using the Contriever. using Contriever-MSMARCO, and input them together with an LM output that may include hallucinations. The model identifies hallucinations, marks phrases or sentences for deletion, and suggests edits.
@@ 6.1 Experiments for Hallucination Detection
Evaluation data.  We use our new benchmark, consisting of 902 annotated passages, [^12] We collected 1010 annotated passages including our experimental batches, excluding earlier batches for evaluations. as our test data. We measure the models’ detection performance based on sentence-level per-category classification task as formulated in Section 3.2. We evaluate both fine-grained detection and binary detection settings. Note that fine-grained hallucination detection is our new proposed task, as prior work mostly focuses on the binary detection setting.
Baselines. Fine-grained hallucination is a new task and we test three baselines that use state-of-the-art proprietary LMs: ChatGPT prompts ChatGPT (gpt-3.5-turbo-0301) with a carefully designed prompt describing all six categories with two demonstrations. GPT4 (gpt-4) does the exact same but prompts GPT4 instead. We were unable to include GPT4 with retrieval in our baseline due to high costs. Since Fava uses retrieval augmentation, for comparability we include Rt+ChatGPT. This baseline prompts ChatGPT using the same prompt and demonstrations but also includes the top five retrieved documents by Contriever at test time to augment the original prompt. [^13] While we also tested strong white box LMs including Llama2-Chat 13B, we found their predictions often show confusion among different categories or violate formats. We additionally evaluate FActScore (Min et al., 2023) for the binary detection task only. This model-based metric prompts ChatGPT and GPT 3 (davinci-003) to decompose a response into a set of atomic facts and verify factuality for each using passages from a designated Wikipedia article.
@@ 6.2 Experiments for Hallucination Editing
Evaluation data.  For editing evaluations, we use open-ended text-generation data and evaluate models’ factuality based on a metric designed to measure the factuality for the target task. In particular, we use the biography generation task proposed by FActScore (Min et al., 2023) to evaluate the effectiveness of editing to reduce hallucinations based on edited outputs’ FActScore results.
@@ 6.3 Human Evaluations
Our automatic evaluations may not fully capture the models’ abilities to detect and refine hallucinations, due to the potential subjectivity of our annotations or the performance of the factuality evaluation metric. We evaluate randomly sampled 50 outputs from Fava as well as the baseline with the highest automatic evaluation score, namely Rt-ChatGPT. We ask human annotators to verify how many of the detection and edits are indeed correct based on the provided retrieved documents. This is similar to our automatic precision evaluation, but instead of coarsely evaluating detection performance at the sentence level, we evaluate the model performance at each detection level.
@@ 7 Results and Analysis
Generator: ChatGPT Generator: Llama2-Chat 70B
Editor  ent  rel  con  inv  subj  unv Avg.  ent  rel  con  inv  subj  unv Avg.
ChatGPT 19.5 28.6 40.0 11.8 7.7 0.0 18.8 24.7 15.6 26.7 11.0 17.6 12.8 24.1
Rt+ChatGPT 28.1 19.2 25.5 5.4 37.7 15.5 24.4 33.7 24.2 24.0 22.2 17.8 4.7 27.8
GPT4 38.6 16.6 17.9 22.2 50.0 17.2 34.2 55.5 60.0 21.2 15.4 2.0 25.0 42.5
Fava (ours) 54.5 25.0 66.7 16.7 70.5 35.3 48.1 57.3 34.5 27.7 52.2 31.3 43.4 47.2
Table 2: Results on fine-grained detection (metric: F1). Full results are in Appendix.
@@ 7.1 Results
Model Chat Llama
Chat 50.1 68.4
R+Chat 64.8 72.8
GPT4 60.8 74.2
FAct 67.7 71.2
Fava 79.6 80.3
Table 3: Results of binary detection.
Generator
Editor CGPT Al-13B Al-7B
@@
Table 4: Results of editing.
@@
Table 5: Ablations of
retrieval for editing.
Fing-grained detection results.  Table 2 shows the fine-grained detection accuracy, namely F1 scores, and binary prediction F1 of Fava and baselines. We provide the full precision and recall scores in Appendix Tables 13 and 14, respectively. Fava significantly outperforms ChatGPT, Ret-ChatGPT, or GPT-4 on both fine-grained error detection and binary error detection. Fava shows high accuracy on error types such as   Sentence ,  Subjective ,  Entity . On the other hand, its performance on  Invented and  Unverifiable is still limited. Those two error types often require intensive search over many documents beyond the top few ones, while Fava by default only considers the top five documents. Table 7.1 shows the binary hallucination detection performance of Fava, baselines models, and FActScore (FAct). Fava outperforms all other models in detecting hallucinations by a large margin.
Editing results.  Table 7.1 shows the results of the editing task on the biography generation task. Our Fava significantly outperforms prompted ChatGPT or Llama2-Chat 70B, despite being a much smaller model in size. We found that ChatGPT is often confused with different errors and tends to over-predict contradictory or subjective statements. Fava shows the largest gains in FActScore, showing the effectiveness of editing and error type detection.
7.2 Analysis
Effects of data scale.  We assess our model on varying sizes of synthetic training data (10k, 20k, and 30k instances) and examine their performance on fine-grained hallucination detection. Appendix Figure 7 demonstrates that Fava variants with a larger number of training instances perform significantly better at detecting fine-grained errors.
Effect of retrieval.  We assess Fava’s editing performance on Alpaca 13B outputs with varying retrieved evidence: top 1 document; top 5 documents; reordering the top 5 documents; top 4 documents and mix them with an introductory paragraph of the target entity (entity matching). Table 5 shows the experimental results. More detailed results are in Appendix Table  18. We found that retrieving the top five documents significantly enhances performance compared to retrieval of only the top one document by 4.5%. Ordering passages using a reranking model gives some improvements, which indicates the non-negligible effect of context order, as reported by prior work (Liu et al., 2023). Including the reference based on entity matching gives a notable enhancement to 50.1%, marking a 3.1% gain from the top five retrievals and 7.6% improvement over the unedited generations. These results indicate although Fava demonstrates strong capabilities, there’s room for improvement in retrieving and incorporating many references.